# SEA: Semantic Map Prediction for Active Exploration of Uncertain Areas

### v1

@@ I INTRODUCTION
With the advancement of embodied agents, research on active exploration in unknown environments has grown rapidly. Exploration tasks in embodied intelligence differ in focus, including discovering and cataloging objects [1, 2], locating specific targets [3, 4, 5], maximizing explored area within limited time [6, 7, 8], and achieving accurate environment reconstruction.
Traditional exploration algorithms rely on hand-crafted optimization methods to maximize coverage. Graph-based methods [9] are precise but inflexible, often falling into local optima, while sampling-based methods [10, 11, 12] handle high-dimensional spaces but cannot guarantee optimality. These approaches generally face a trade-off between global awareness and exploration efficiency. To overcome this, reinforcement learning (RL) and deep learning (DL) have been introduced, leveraging neural networks’ implicit long-term memory to infer global structure [13]. However, studies on spatial intelligence reveal that neural networks often fail to capture topology and lack robust 3D perception [14], limiting their reliability in complex scenes.
@@
To address these limitations, we integrate the adaptive planning capabilities of RL with explicit spatial reasoning to achieve more efficient and flexible exploration. Inspired by vision-and-language navigation (VLN), where semantic maps improve memory [15, 16], we argue that learning-based exploration should return to its core goal: actively and comprehensively covering unknown environments while building complete and accurate maps for decision-making.
@@ II-A Semantic Scene Completion
A concept closely related to our work is Semantic Scene Completion (SSC), which aims to infer the complete geometry and semantics of a scene from sparse inputs like depth maps or monocular RGB images. Early work primarily focused on object-level completion, reconstructing full geometry from limited 2D/3D observations [17, 18]. With the rise of deep learning, research has shifted toward completing full semantic scenes by jointly predicting semantics and geometry [19, 20, 21]. However, a key distinction from our approach is that most SSC methods reconstruct only the current visible scene in dense 3D space, without predicting unseen areas beyond the field of view.
Several studies combine completion with reconstruction and exploration [22, 23, 24]. For instance, PredRecon [22] applies object-level completion to guide UAV waypoint selection for reconstruction, yet it focuses on object geometry rather than scene-level exploration. The heuristic-based method in [24] identifies unknown frontiers and applies rule-based strategies to guide exploration. Although it integrates unknown-area prediction, it only distinguishes between explored and unexplored regions and lacks perceptual priors for active exploration.
@@ II-B Occupancy Mapping and Semantic Mapping
The objective of mapping is to construct a structured representation of an environment. Depending on whether semantic information is considered, this can be categorized into occupancy and semantic mapping. In occupancy mapping, Chaplot et al. introduced an active navigation framework that projects egocentric views to generate obstacle maps [6]. Other works construct obstacle maps [25, 26, 2, 27] or build topological maps capturing spatial relations among rooms and objects [28, 29, 30, 31, 32, 33, 34].
While some works [35, 36] attempt to enhance environmental observations by optimizing camera rotation for next-best-view selection, they do not guide the agent toward active exploration for improved semantic mapping. Similarly, studies on optimizing exploration perspectives for point cloud reconstruction [37, 38, 39, 40] adopt a next-best-view approach but are constrained by the high storage cost and computational resources required for 3D point clouds. As a result, many semantic mapping approaches struggle to integrate with autonomous exploration tasks.
@@ II-C Active Exploration
Most existing exploration approaches [6, 41, 8, 7, 26, 1, 2] focus on maximizing coverage or object diversity but lack explicit semantic reasoning. The works most related to ours are SSCNav[42] and L2M [43], which leverage predicted semantic local maps and uncertainty estimation. However, they are limited to object-goal navigation. SSCNav discards local maps at each step, preventing global map accumulation, and its confidence maps are used to correct errors rather than guide exploration in uncertain regions. L2M also exploits uncertainty, but only for specific object categories relevant to its navigation task, without extending to global semantic completion.
The recent work GLEAM [44] focuses on large-scale training and cross-dataset policy generalization. Its focus on scalability is complementary to ours; while GLEAM advances generalizable exploration, our method addresses the critical need for semantic uncertainty and completion to construct high-fidelity maps under a limited budget. To our knowledge, no prior work simultaneously explores unseen indoor environments and constructs predictive semantic maps to guide the process. We argue that semantic reasoning, fundamental to how humans explore, is a highly promising direction for enhancing both efficiency and scene understanding in embodied AI.
@@ III-B ASC-based Local Mapper
Building a complete map to represent the global environment requires fusing information from egocentric views. Existing DRL-based approaches [45] directly accumulate maps by projecting only the areas visible to the agent, leading to inefficient mapping. We highlight that, with prior knowledge of indoor scenes, unseen areas can be predicted and completed, enhancing the semantic map with broader and more accurate coverage.
@@ III-B1 Panoramic Segmented Local Map
In this module, we first perform panoramic segmentation of RGB observation according to the 40 semantic categories of the MP3D dataset. Our panoramic semantic segmentation employs off-the-shelf ACNet [46] and leverages the pretrained panoramic segmentation model from SSCNav [42] which is trained with 209,200 RGB-D images of 40 object categories from MP3D training houses. As a comparison, SemExp [4] utilizes only 16 semantic categories relevant to the Habitat Object Goal Navigation task. In contrast, our local maps include all 40 semantic categories plus an additional void category. Moreover, our method provides a panoramic semantic view, excluding unseen areas such as floors and walls.
@@
We emphasize that, although we do not explore real-world data expansion in this paper, we have aimed to replicate real-world noise conditions in this section to facilitate future research. To align with the semantic noise encountered in real-world applications, we employ a segmentation model for semantic segmentation instead of using the ground-truth annotations. Regarding pose estimation, while we obtain ground-truth poses directly from the simulator, this process can be replicated in the real world using a SLAM front end. Considering that common front-end methods like VINS can achieve an accuracy of 7 cm ATE in a 147 m trajectory[47], we believe this assumption is reasonable.
@@ III-B2 Local Map Completion
As illustrated in Figure 3, we calculate the unseen area in $m_{proj}$ to obtain a mask $mask_{us}\in\{0,1\}^{M\times M}$. We exploit ResNet with a fully convolutional neural network with 1 maxpooling layer, 4 down-sample residual blocks, and 5 up-sample residual blocks, following the same architecture in SSCNav [48]. Differently, we use only one observed projected map $m_{proj}$ to get the predicted map $m_{pred}$ of the same size, and fill the unseen part of $m_{proj}$ with that in $m_{pred}$ to form $m_{cmplt}$ for consistency. The loss is calculated as the Cross Entropy Loss between $m_{cmplt}$ and $m_{gt}$.
@@ III-C1 Confidence-aware Long-term Policy
Existing approaches [6, 1] select long-term goals across the entire map. However, our experiments show that these methods often select goals near map corners, limiting their ability to utilize environmental information effectively. This inefficiency negatively impacts the overall exploration process.
To address these limitations, we propose leveraging the Soft Actor-Critic (SAC) algorithm [49], an off-policy, model-free reinforcement learning method that enables more effective long-term goal selection. Within a single long-term goal period, the agent’s reachable distance remains limited to the $M\times M$} local map. This allows us to select a new long-term goal based on the current $m_{proj}$, $m_{cmplt}$ and $m_{conf}$. The interval for long-term goal planning is set to $\mu=25$. The outputs of the Local Mapper, including $m_{proj}$, $m_{cmplt}$ and $m_{conf}$, are fed into the SAC module for exploration. Additionally, the embeddings of the agent’s current position and orientation are included as inputs.
@@ III-C2 Short-term Policy
For the short-term policy, a deterministic path planning method, Fast Marching Method (FMM) [50], is used. Based on $m_{proj}$, FMM plans a path toward the selected long-term goal and determines the agent’s next action. It maintains an obstacle local map, allowing the agent to navigate efficiently while avoiding obstacles during exploration.
@@ IV-A Experiment Setup
Dataset and Simulator. We use Habitat [48] as the simulation platform and MP3D [51] as the dataset. MP3D is an RGB-D indoor dataset with 90 large-scale scenes: 61 for training, 11 for validation, and 18 for testing. Our method is trained with 1M observation frames from the training split and evaluated with 50k frames from the validation split. For semantic completion and confidence modeling, we adopt the local map completion dataset from SSCNav, which contains over 52.3k pairs of ground-truth and projected maps. Ground-truth full maps $M_{gt}$ for evaluation are generated following SMNet [52] using MP3D semantics.
@@
Following [7], the maximum exploration step is set to 500. The local map resolution is $128\times 128$, and the full map is $1280\times 1280$ (corresponding to $48\times 48,m^{2}$ in the physical world). Although insufficient for full exploration, this effectively highlights differences across strategies for semantic map construction. Long-term goals are updated every 10 steps. Empirically, ASC is about half of total coverage. We estimate $\Delta_{asc}$ by doubling it and compute $\Delta_{conf}$ by scaling $\Delta_{asc}$ by 1000. The final reward (Eq. 2) is calculated with hyperparameters $\eta_{c}=0.6$, $\eta_{asc}=0.4$, and $\eta_{conf}=0.4$.
@@ IV-B Baselines
  • SemExp* [4]: Originally designed for object-goal navigation, SemExp shares its architecture with ANS[6], which targets occupied coverage. We modify its reward to prioritize coverage and retrain it on the same dataset, effectively converting it into an exploration method.
  • Impact [7]: Trained with purely intrinsic rewards, Impact explores to maximize occupied coverage. We evaluate it using the provided trained model under the same setting reported in the original paper.
  • EE [2]: This work analyzes how reward design affects coverage and object detection. We adopt the released model with the best coverage performance for evaluation.
  • ARiADNE [53] is a state-of-the-art RL-based exploration method that outperforms classical approaches like TARE [12] on grid maps. For a fair comparison, we adapted it to match our method’s mapping scope and integrated our short-term policy, FMM, as its fine-grained exploration strategy in the continuous Habitat environment.
@@ IV-C Results on MP3D Datasets
Notably, the best coverage reported in Impact [7] when evaluated on the MP3D Val split is 144.64 $m^{2}$. However, this difference arises because our experimental setup limits the depth range of map projection to 5 m, whereas the mapper in Impact’s code uses 10 m. Consequently, the coverage of our semantic map mapper appears lower than in Impact’s original paper. Nevertheless, since all experiments are conducted under a consistent setting in our study, the results remain valid and reliable.

### v2

@@ I INTRODUCTION
With the advancement of embodied agents, research on active exploration in unknown environments has grown rapidly. Exploration tasks in embodied intelligence differ in focus, including discovering and cataloging objects [chaplot2020semantic, ramakrishnan2021exploration], locating specific targets [ye2021auxiliaryob, chaplot2020object, zhu2022navigating], maximizing explored area within limited time [chaplot2020learning, bigazzi2022focus, juncheng2021learning], and achieving accurate environment reconstruction.
Traditional exploration algorithms rely on hand-crafted optimization methods to maximize coverage. Graph-based methods [dolgov2008practical] are precise but inflexible, often falling into local optima, while sampling-based methods [zhou2019robust, zhou2021fuel, Cao2021TAREAH] handle high-dimensional spaces but cannot guarantee optimality. These approaches generally face a trade-off between global awareness and exploration efficiency. To overcome this, reinforcement learning (RL) and deep learning (DL) have been introduced, leveraging neural networks’ implicit long-term memory to infer global structure [cao2024dare]. However, studies on spatial intelligence reveal that neural networks often fail to capture topology and lack robust 3D perception [yang2024thinking], limiting their reliability in complex scenes.
@@
To address these limitations, we integrate the adaptive planning capabilities of RL with explicit spatial reasoning to achieve more efficient and flexible exploration. Inspired by vision-and-language navigation (VLN), where semantic maps improve memory [Wang2023GridMMGM, Chen2022ThinkGA], we argue that learning-based exploration should return to its core goal: actively and comprehensively covering unknown environments while building complete and accurate maps for decision-making.
@@ II-A Semantic Scene Completion
A concept closely related to our work is Semantic Scene Completion (SSC), which aims to infer the complete geometry and semantics of a scene from sparse inputs like depth maps or monocular RGB images. Early work primarily focused on object-level completion, reconstructing full geometry from limited 2D/3D observations [Yuan2018PCNPC, Tchapmi2019TopNetSP]. With the rise of deep learning, research has shifted toward completing full semantic scenes by jointly predicting semantics and geometry [Song2016SemanticSC, Li2023VoxFormerSV, jiang2023symphonies]. However, a key distinction from our approach is that most SSC methods reconstruct only the current visible scene in dense 3D space, without predicting unseen areas beyond the field of view.
Several studies combine completion with reconstruction and exploration [feng2023predrecon, Gao2024ActiveVP, Xu2023HeuristicbasedIP]. For instance, PredRecon [feng2023predrecon] applies object-level completion to guide UAV waypoint selection for reconstruction, yet it focuses on object geometry rather than scene-level exploration. The heuristic-based method in [Xu2023HeuristicbasedIP] identifies unknown frontiers and applies rule-based strategies to guide exploration. Although it integrates unknown-area prediction, it only distinguishes between explored and unexplored regions and lacks perceptual priors for active exploration.
@@ II-B Occupancy Mapping and Semantic Mapping
The objective of mapping is to construct a structured representation of an environment. Depending on whether semantic information is considered, this can be categorized into occupancy and semantic mapping. In occupancy mapping, Chaplot et al. introduced an active navigation framework that projects egocentric views to generate obstacle maps [chaplot2020learning]. Other works construct obstacle maps [chen2019learning, ramakrishnan2020occupancy, ramakrishnan2021exploration, sharma2022occupancy] or build topological maps capturing spatial relations among rooms and objects [morad2021graph, wang2022lightweight, li2022remote, dang2022unbiased, chaplot2020neural, ravichandran2022hierarchical, zheng2021research].
While some works [jayaraman2016look, jayaraman2018learning] attempt to enhance environmental observations by optimizing camera rotation for next-best-view selection, they do not guide the agent toward active exploration for improved semantic mapping. Similarly, studies on optimizing exploration perspectives for point cloud reconstruction [guedon2023macarons, cieslewski2017rapid, vasquez2014volumetric, vasquez2017view] adopt a next-best-view approach but are constrained by the high storage cost and computational resources required for 3D point clouds. As a result, many semantic mapping approaches struggle to integrate with autonomous exploration tasks.
@@ II-C Active Exploration
Most existing exploration approaches [chaplot2020learning, liu2021symmetry, juncheng2021learning, bigazzi2022focus, ramakrishnan2020occupancy, chaplot2020semantic, ramakrishnan2021exploration] focus on maximizing coverage or object diversity but lack explicit semantic reasoning. The works most related to ours are SSCNav[liang2021sscnav] and L2M [georgakis2021learning], which leverage predicted semantic local maps and uncertainty estimation. However, they are limited to object-goal navigation. SSCNav discards local maps at each step, preventing global map accumulation, and its confidence maps are used to correct errors rather than guide exploration in uncertain regions. L2M also exploits uncertainty, but only for specific object categories relevant to its navigation task, without extending to global semantic completion.
The recent work GLEAM [chen2025gleam] focuses on large-scale training and cross-dataset policy generalization. Its focus on scalability is complementary to ours; while GLEAM advances generalizable exploration, our method addresses the critical need for semantic uncertainty and completion to construct high-fidelity maps under a limited budget. To our knowledge, no prior work simultaneously explores unseen indoor environments and constructs predictive semantic maps to guide the process. We argue that semantic reasoning, fundamental to how humans explore, is a highly promising direction for enhancing both efficiency and scene understanding in embodied AI.
@@ III-B ASC-based Local Mapper
Building a complete map to represent the global environment requires fusing information from egocentric views. Existing DRL-based approaches [li2023autonomous] directly accumulate maps by projecting only the areas visible to the agent, leading to inefficient mapping. We highlight that, with prior knowledge of indoor scenes, unseen areas can be predicted and completed, enhancing the semantic map with broader and more accurate coverage.
@@ III-B1 Panoramic Segmented Local Map
In this module, we first perform panoramic segmentation of RGB observation according to the 40 semantic categories of the MP3D dataset. Our panoramic semantic segmentation employs off-the-shelf ACNet [hu2019acnet] and leverages the pretrained panoramic segmentation model from SSCNav [liang2021sscnav] which is trained with 209,200 RGB-D images of 40 object categories from MP3D training houses. As a comparison, SemExp [chaplot2020object] utilizes only 16 semantic categories relevant to the Habitat Object Goal Navigation task. In contrast, our local maps include all 40 semantic categories plus an additional void category. Moreover, our method provides a panoramic semantic view, excluding unseen areas such as floors and walls.
@@
We emphasize that, although we do not explore real-world data expansion in this paper, we have aimed to replicate real-world noise conditions in this section to facilitate future research. To align with the semantic noise encountered in real-world applications, we employ a segmentation model for semantic segmentation instead of using the ground-truth annotations. Regarding pose estimation, while we obtain ground-truth poses directly from the simulator, this process can be replicated in the real world using a SLAM front end. Considering that common front-end methods like VINS can achieve an accuracy of 7 cm ATE in a 147 m trajectory[Campos2020ORBSLAM3AA], we believe this assumption is reasonable.
@@ III-B2 Local Map Completion
As illustrated in Figure 3, we calculate the unseen area in $m_{proj}$ to obtain a mask $mask_{us}\in\{0,1\}^{M\times M}$. We exploit ResNet with a fully convolutional neural network with 1 maxpooling layer, 4 down-sample residual blocks, and 5 up-sample residual blocks, following the same architecture in SSCNav [savva2019habitat]. Differently, we use only one observed projected map $m_{proj}$ to get the predicted map $m_{pred}$ of the same size, and fill the unseen part of $m_{proj}$ with that in $m_{pred}$ to form $m_{cmplt}$ for consistency. The loss is calculated as the Cross Entropy Loss between $m_{cmplt}$ and $m_{gt}$.
@@ III-C1 Confidence-aware Long-term Policy
Existing approaches [chaplot2020learning, chaplot2020semantic] select long-term goals across the entire map. However, our experiments show that these methods often select goals near map corners, limiting their ability to utilize environmental information effectively. This inefficiency negatively impacts the overall exploration process.
To address these limitations, we propose leveraging the Soft Actor-Critic (SAC) algorithm [haarnoja2018soft], an off-policy, model-free reinforcement learning method that enables more effective long-term goal selection. Within a single long-term goal period, the agent’s reachable distance remains limited to the $M\times M$} local map. This allows us to select a new long-term goal based on the current $m_{proj}$, $m_{cmplt}$ and $m_{conf}$. The interval for long-term goal planning is set to $\mu=25$. The outputs of the Local Mapper, including $m_{proj}$, $m_{cmplt}$ and $m_{conf}$, are fed into the SAC module for exploration. Additionally, the embeddings of the agent’s current position and orientation are included as inputs.
@@ III-C2 Short-term Policy
For the short-term policy, a deterministic path planning method, Fast Marching Method (FMM) [sethian1996fast], is used. Based on $m_{proj}$, FMM plans a path toward the selected long-term goal and determines the agent’s next action. It maintains an obstacle local map, allowing the agent to navigate efficiently while avoiding obstacles during exploration.
@@ IV-A Experiment Setup
Dataset and Simulator. We use Habitat [savva2019habitat] as the simulation platform and MP3D [chang2017matterport3d] as the dataset. MP3D is an RGB-D indoor dataset with 90 large-scale scenes: 61 for training, 11 for validation, and 18 for testing. Our method is trained with 1M observation frames from the training split and evaluated with 50k frames from the validation split. For semantic completion and confidence modeling, we adopt the local map completion dataset from SSCNav, which contains over 52.3k pairs of ground-truth and projected maps. Ground-truth full maps $M_{gt}$ for evaluation are generated following SMNet [cartillier2021semantic] using MP3D semantics.
@@
Following [bigazzi2022focus], the maximum exploration step is set to 500. The local map resolution is $128\times 128$, and the full map is $1280\times 1280$ (corresponding to $48\times 48,m^{2}$ in the physical world). Although insufficient for full exploration, this effectively highlights differences across strategies for semantic map construction. Long-term goals are updated every 10 steps. Empirically, ASC is about half of total coverage. We estimate $\Delta_{asc}$ by doubling it and compute $\Delta_{conf}$ by scaling $\Delta_{asc}$ by 1000. The final reward (Eq. 2) is calculated with hyperparameters $\eta_{c}=0.6$, $\eta_{asc}=0.4$, and $\eta_{conf}=0.4$.
@@ IV-B Baselines
  • SemExp* [chaplot2020object]: Originally designed for object-goal navigation, SemExp shares its architecture with ANS[chaplot2020learning], which targets occupied coverage. We modify its reward to prioritize coverage and retrain it on the same dataset, effectively converting it into an exploration method.
  • Impact [bigazzi2022focus]: Trained with purely intrinsic rewards, Impact explores to maximize occupied coverage. We evaluate it using the provided trained model under the same setting reported in the original paper.
  • EE [ramakrishnan2021exploration]: This work analyzes how reward design affects coverage and object detection. We adopt the released model with the best coverage performance for evaluation.
  • ARiADNE [Cao2023ARiADNEAR] is a state-of-the-art RL-based exploration method that outperforms classical approaches like TARE [Cao2021TAREAH] on grid maps. For a fair comparison, we adapted it to match our method’s mapping scope and integrated our short-term policy, FMM, as its fine-grained exploration strategy in the continuous Habitat environment.
@@ IV-C Results on MP3D Datasets
Notably, the best coverage reported in Impact [bigazzi2022focus] when evaluated on the MP3D Val split is 144.64 $m^{2}$. However, this difference arises because our experimental setup limits the depth range of map projection to 5 m, whereas the mapper in Impact’s code uses 10 m. Consequently, the coverage of our semantic map mapper appears lower than in Impact’s original paper. Nevertheless, since all experiments are conducted under a consistent setting in our study, the results remain valid and reliable.